OpenStudy (anonymous):

Let z = cos((2pi)/n) + i sin((2pi)/n) and n ≥ 2. Show that 1 + z + · · · + z^(n−1) = 0.

hartnn (hartnn):

can you find z^n ?

hartnn (hartnn):

then we use the formula \(\large z^n-1= (z-1)(1+z+z^2+...+z^{n-1}) \\ z \ne1\)

hartnn (hartnn):

see whether you get z^n=1 or not..

OpenStudy (anonymous):

z^n can be found using deMoivre's theorem, in which case z^n would become cos((2pi)) + i sin((2pi)) = 1.

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

and then from the formula that you said it shows that it is 0. Thanks.

hartnn (hartnn):

welcome ^_^